### v1

# GlobalGeoTree: A Multi-Granular Vision-Language Dataset for Global Tree Species Classification

Abstract
Global tree species mapping using remote sensing data is vital for biodiversity monitoring, forest management, and ecological research. However, progress in this field has been constrained by the scarcity of large-scale, labeled datasets. To address this, we introduce GlobalGeoTree—a comprehensive global dataset for tree species classification. GlobalGeoTree comprises 6.3 million geolocated tree occurrences, spanning 275 families, 2,734 genera, and 21,001 species across the hierarchical taxonomic levels. Each sample is paired with Sentinel-2 image time series and 27 auxiliary environmental variables, encompassing bioclimatic, geographic, and soil data. The dataset is partitioned into GlobalGeoTree-6M for model pretraining and curated evaluation subsets, primarily GlobalGeoTree-10kEval for zero-shot and few-shot benchmarking. To demonstrate the utility of the dataset, we introduce a baseline model, GeoTreeCLIP, which leverages paired remote sensing data and taxonomic text labels within a vision-language framework pretrained on GlobalGeoTree-6M. Experimental results show that GeoTreeCLIP achieves substantial improvements in zero- and few-shot classification on GlobalGeoTree-10kEval over existing advanced models. By making the dataset, models, and code publicly available, we aim to establish a benchmark to advance tree species classification and foster innovation in biodiversity research and ecological applications.
Figure 1: Overview of the GlobalGeoTree dataset, which includes 6.3 million samples spanning 21,001 tree species across 221 countries/regions. The map illustrates the geographic coverage, with color intensity representing the number of samples in each 1° × 1° latitude/longitude grid. Each sample is paired with remote sensing data, including Sentinel-2 time series, auxiliary environmental variables, and hierarchical taxonomic labels spanning from functional type to species level.
1 Introduction
Forests cover approximately 31% of the global land surface [1] and provide essential ecosystem services, including carbon sequestration [2], biodiversity conservation [3], and climate regulation [4]. Accurate and large-scale mapping of tree species plays an increasingly vital role in addressing pressing environmental challenges [5], including effective biodiversity monitoring [6], informed forest management practices [7], and comprehensive ecological research aimed at understanding the complex impacts of climate change [8].
Traditional ground-based forest monitoring methods [9], while providing detailed information, are often limited in their spatial and temporal coverage, making it challenging to obtain a comprehensive understanding of global forest composition and dynamics. In contrast, remote sensing has emerged as a key technology for large-scale forest monitoring, offering non-invasive and cost-effective approaches to tree species classification [10]. Despite significant advancements in this field, progress has been constrained by the limited availability of comprehensive, high-quality, and accurately labeled datasets that capture the global diversity of tree species [11]. Existing datasets typically focus on specific geographic regions or limited taxonomic coverage, hampering the development of models with global applicability [12].
To bridge these gaps, we present GlobalGeoTree, a large-scale dataset comprising 6.3 million remote sensing samples paired with multi-level taxonomic labels. This dataset integrates time-series satellite imagery from Sentinel-2 with 27 bioclimatic, geographic, and soil variables, offering a rich multimodal representation of tree species within their environmental contexts. The taxonomic hierarchy spans family, genus, and species levels, enabling classification across various scales of biological organization.
In addition to the dataset, we introduce GeoTreeCLIP, a vision-language model specifically designed for tree species classification. Drawing on frameworks like CLIP [13], our approach aligns satellite imagery with taxonomic labels to learn nuanced representations. Unlike traditional classifiers treating labels as discrete categories, vision-language models can inherently process label hierarchical structure, generalizing to unseen species through representations of related genera or families [14]. This enables robust zero-shot and few-shot learning of GeoTreeCLIP, which are critical for addressing the vast scale of global biodiversity, ever-evolving species catalogs, and the practical impossibility of exhaustive data collection for all taxa.
GeoTreeCLIP leverages domain-specific pretraining on GlobalGeoTree-6M, main part of the dataset tailored for model pretraining, and evaluated on a specialized benchmark, GlobalGeoTree-10kEval, which enables a comprehensive assessment of model performance across multiple taxonomic levels. Through the open availability of GlobalGeoTree, its associated models, and evaluation protocols, we seek to establish a community-driven benchmarking standard that will accelerate the development of generalizable models for tree species mapping and deepen our understanding of global forest biodiversity.
2 Related work
2.1 Open datasets for tree species classification
Table 1 provides an overview of notable open datasets for tree species classification, detailing their geographic coverage, size, taxonomic diversity, and publication year. Several datasets have contributed to tree species classification. For instance, the Seu Nico Forest dataset [15] from Brazil provides geolocated samples for 228 species but is geographically constrained. Similarly, the Maraca Ecological Station dataset [16] includes 110 species but is also region-specific. In Europe, the EUForest dataset [17] offers broader coverage with data for 242 species. On a global scale, datasets such as Tallo [18] provide significant taxonomic diversity, covering 5,163 species across 187 families. However, these datasets lack integration with remote sensing or environmental variables, limiting their application in ecological modeling.
Table 1: Overview of publicly available datasets for tree species classification.
Dataset Geographic scope Size Classes Year
Seu Nico Forest [15] Brazil 2,868 54 families; 139 genera; 228 species 2015
EUForest [17] Europe 588,983 83 genera; 242 species 2017
Maraca Eco. Sta. [16] Brazil 680 40 families; 110 species 2020
TreeSatAI [19] Germany 50,381 15 genera; 20 species 2022
Tallo [18] Global 498,839 187 families; 1,453 genera; 5,163 species 2022
Indi. Tree Point Clouds [20] Germany 1,491 22 species 2022
NEON Veg. Struc. [21] USA N/A 949 genera; 2,826 species 2023
PureForest [22] France 135,569 18 species 2024
Planted [23] Global 2,264,747 46 genera; 40 species 2024
GlobalGeoTree Global 6,263,345 275 families; 2,734 genera; 21,001 species 2025
Advances in high-resolution imaging and lidar technologies have enabled datasets like PureForest [22] and Individual Tree Point Clouds [20], which utilize aerial and point cloud data for species classification. While these datasets offer detailed structural information, they remain region-specific and lack the spectral and temporal depth of satellite-based datasets. The TreeSatAI dataset [19] combines multi-sensor data, including aerial imagery and Sentinel-1/2, for tree species classification in Germany but covers only 20 species. Similarly, the Planted dataset [23] focuses on only 40 planted species globally, limiting its broader applicability.
While valuable, these existing datasets highlight the persistent need for a benchmark that synergizes global coverage, deep taxonomic information, and multimodal remote sensing data, a gap GlobalGeoTree aims to fill.
2.2 Vision-language models for remote sensing applications
Vision-language models (VLMs) enable the integration of visual and textual information. Among these, Contrastive Language-Image Pretraining (CLIP) [13] has demonstrated exceptional zero-shot transfer capabilities by jointly training image and text encoders through a contrastive learning objective, aligning image-text pairs within a shared embedding space. For tree species classification, CLIP’s ability to learn from image-text pairings (such as satellite imagery and taxonomic labels) offers a path to capture complex visual and semantic relationships. Its proven zero-shot capabilities are particularly suited for addressing the challenges of identifying species within dynamic and evolving catalogs [14]. Meanwhile, its few-shot capabilities tackle the issue of limited labeled data, a common obstacle in biodiversity research, offering an advantage over traditional supervised methods.
In remote sensing, VLMs have been applied to various tasks, including image classification, retrieval, and scene understanding, with domain-specific adaptations yielding significant improvements. For example, RemoteCLIP [24], the first VLM specifically tailored for remote sensing, leverages pretraining on large-scale remote sensing imagery paired with aligned text, achieving state-of-the-art performance in zero-shot classification, linear probing, and few-shot learning. Similarly, SkyCLIP [25] and GeoLangBind [26] extend the capabilities of CLIP through continual pretraining on semantically diverse remote sensing image-text pairs. These models demonstrate enhanced generalization and transferability, achieving substantial gains in tasks such as zero-shot scene classification, fine-grained classification, and cross-modal retrieval compared to the original CLIP model.
These advancements underscore the importance of domain-specific pretraining in adapting VLMs for remote sensing applications. Aligning models more closely with the unique characteristics of remote sensing tasks has demonstrated significant potential to advance progress in this field.
3 The GlobalGeoTree dataset
3.1 Geolocated data collection and preprocessing
The GlobalGeoTree dataset provides unprecedented global and taxonomic coverage for tree species classification using remote sensing data, and the collection involved several key steps:
Tree species catalog construction
We constructed a comprehensive tree species catalog by integrating two major global repositories: TreeGOER [27] and GlobalTreeSearch [28], containing 48,129 and 57,681 tree species respectively. This compilation was further enriched with multiple open-source datasets documented in Table 1. The taxonomic framework was subsequently validated and expanded using the Global Biodiversity Information Facility (GBIF) Species API [29], ensuring nomenclatural consistency and accuracy. The resulting catalog encompasses 87,845 species, representing the global diversity of tree species.
Geolocation sampling
For each tree species in our catalog, we queried the GBIF Occurrence API [30] to retrieve global geolocations with documented occurrences. To ensure data quality and reliability, we applied strict filtering criteria, including: (1) selecting only recent observations recorded between 2015 and 2024; (2) limiting data to human observation records; (3) excluding records with geospatial issues as flagged by GBIF (e.g., country-coordinate mismatches); (4) filtering for occurrences with a "present" status; and (5) removing duplicate entries and observations with low geographic precision. Additionally, we ensured all samples conform to open data licenses (CC0 1.0, CC-BY-4.0, etc.), maintaining the dataset’s accessibility and reusability for the broader research community.
Forest layer filtering
To ensure that our dataset focuses on actual forest areas rather than street trees, parks, or urban vegetation, we utilized the EC JRC Global Map of Forest Cover 2020 [31] at 10m resolution. This filtering step eliminated potential samples from non-forest environments, improving the dataset’s relevance for forestry and ecological applications.
3.2 Paired remote sensing data
The resulting GlobalGeoTree dataset comprises 6,263,345 samples distributed across 221 countries and regions. More details can be checked in Appendix A. Table 2 provides an overview of the features included in each sample, which consist of paired Earth Observation (EO) data and auxiliary environmental variables derived from remote sensing sources.
Table 2: Overview of features in each sample in the GlobalGeoTree dataset.
Feature Name Type Description
Remote Sensing Data
Sentinel-2 Time Series float 12 monthly median composites; Includes RGB, NIR, Vegetation Red Edge, and SWIR bands; dimensions: (12, 10, 5, 5).
Geographic Variables float Latitude and longitude, as well as elevation, slope, and aspect derived from USGS (SRTM) (30m resolution).
Soil Variables float 3 Volumetric Water Content data at 33kPa across depths: 0-5 cm, 5-15 cm, 15-30 cm (250m resolution).
Bioclim Variables float 19 climatic variables from WorldClim (1km resolution).
Text Labels
level0 string Functional type of the species (e.g., Evergreen Broadleaf).
level1_family string Taxonomic family of the species (e.g., Myrtaceae).
level2_genus string Taxonomic genus of the species (e.g., Eucalyptus).
level3_species string Scientific name of the species (e.g., Eucalyptus kybeanensis).
Meta Data
location string Geographic location of the sample (e.g., Australia).
country_code string ISO country code of the sample location.
source string Source of the sample record (e.g., iNaturalist Research-grade Observations).
species_key float Unique identifier for the species in the GBIF database.
record_year int The year when the record was collected.
EO data
The Sentinel-2 data for each sample consists of a time series of 12 monthly median composites from January to December 2020. For each month, all L2A Sentinel-2 images with less than 30% cloud cover were collected, and the median composites were generated from these images. For each composite, a 5×5 pixel patch centered on the geolocation of the tree species is included. The selected patch size accounts for typical crown sizes (10–30 m) [18] and aligns with other public datasets (e.g., PureForest [22], TreeSatAI [19]), providing valuable spatial context around the specific location. The full-year temporal coverage enables models to capture the phenological patterns exhibited by different tree species across seasons [5].
Auxiliary data
The auxiliary data enriches the EO data for each sample by providing additional contextual environmental information. Geographic variables, such as elevation, slope, and aspect, are derived from the USGS SRTM dataset [32], while soil data, including volumetric water content at 33 kPa across three depths (0–5 cm, 5–15 cm, 15–30 cm), are obtained from SoilGrid [33]. Additionally, 19 bioclimatic variables are sourced from WorldClim [34]. Due to the relatively coarse spatial resolution of these datasets (ranging from 30m to 1km), only the values corresponding to the exact coordinates of each occurrence are extracted to ensure precision and relevance [35].
3.3 Dataset partitioning
For effective model development and evaluation, the GlobalGeoTree dataset was partitioned into GlobalGeoTree-6M and curated evaluation subsets, primarily GlobalGeoTree-10kEval.
Figure 2: Geographic distribution of GlobalGeoTree-10kEval. This benchmark includes species selected from Frequent, Common, and Rare categories, as described in the text.
GlobalGeoTree-6M comprises the vast majority of the samples and is specifically designed for model pretraining. This large size allows models to learn robust and generalizable representations of tree species and their associated environmental contexts.
Figure 3: Species in GlobalGeoTree are categorized into Frequent, Common and Rare groups based on the samples per species.
GlobalGeoTree-10kEval, is a carefully curated dataset intended for benchmarking model performance across taxonomic levels and species frequency categories in a fair and robust manner. To address the characteristic long-tail distribution observed in the datasets (detailed in Appendix A.2), species in GlobalGeoTree dataset were categorized into three groups based on the number of samples: Frequent (more than 1500), Common (100–1500), and Rare (less than 100), as shown in Fig. 3.
The GlobalGeoTree-10kEval dataset includes 30 species from each of these three categories, resulting in a total of 90 species. The sample proportions within this evaluation set are 12% for Rare species, 33% for Common species, and 55% for Frequent species, culminating in around 10,000 samples. Fig. 2 shows the geographical distribution of GlobalGeoTree-10kEval, which spans diverse regions across the globe. This global distribution ensures that the dataset captures a wide range of ecological and environmental contexts, making it representative of real-world scenarios. By focusing on a diverse set of species with varying levels of representation, GlobalGeoTree-10kEval serves as a robust evaluation benchmark for assessing the ability of models to tackle challenges posed by the long-tail distribution of species and shifts in geographical domains.
Beyond GlobalGeoTree-10kEval, we developed larger evaluation sets—GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900—by selecting 100 and 300 species from each category, respectively. Crucially, all samples within these evaluation sets, are excluded from the GlobalGeoTree-6M pretraining set to ensure fair evaluation. Details of all evaluation subsets (Appendix B) and the corresponding evaluation results (Appendix C) are also provided. Given the complexity of global tree species classification, our primary analysis focuses on the 90-species GlobalGeoTree-10kEval, which serves as a practical starting point for systematic benchmarking.
4 Benchmarks
4.1 GeoTreeCLIP model
To establish baseline performance on the GlobalGeoTree dataset, we developed GeoTreeCLIP, a vision-language model specifically designed for tree species classification. The core motivation for adopting the CLIP architecture lies in its ability to learn general and powerful representations through multimodal contrastive learning, which jointly trains on image-text pairs at scale. This approach not only aligns visual and textual modalities but also demonstrates strong transfer capabilities, particularly in zero-shot and few-shot scenarios [14]. Such capabilities are especially valuable given the continuously expanding species catalog and the practical limitations of obtaining exhaustive labeled data for all taxa. These qualities make the CLIP architecture particularly suitable for tree species classification, where leveraging rich textual descriptions (e.g., multi-level taxonomic labels) enhances image understanding and addresses challenges such as rare species representation and complex hierarchical structures.
Figure 4: Architecture of the GeoTreeCLIP baseline model. It processes Sentinel-2 time series and auxiliary data through visual encoder and MLP, and hierarchical taxonomic labels through a text encoder. The resulting multimodal visual and text features are then aligned using a contrastive loss.
As shown in Fig. 4, the GeoTreeCLIP model architecture consists of the following components:
  • Visual Encoder: A ViT-B/16 backbone [36] augmented with a temporal attention [37] mechanism to process 12-month Sentinel-2 time series data.
  • Auxiliary Feature Integration: A multi-layer perceptron (MLP) [38, 35] designed to process bioclimatic, soil, and geographic data.
  • Text Encoder: A 77-token causal autoregressive transformer [37, 39] that encodes taxonomic text data across functional type, family, genus, and species levels.
Pretraining details
Both the Visual Encoder and Text Encoder are initialized using OpenAI’s publicly available CLIP checkpoint [13] and further pre-trained on the GlobalGeoTree-6M using Distributed Data Parallel (DDP) across 5 NVIDIA 3090 (24GB) GPUs. We employed a batch size of 384 per GPU, with gradient accumulation over 2 steps, resulting in an effective batch size of 768 per GPU (3840 globally). The AdamW optimizer [40] was used with a base learning rate of $1\times 10^{-5}$ for the visual and auxiliary encoders, and a reduced learning rate of $1\times 10^{-6}$ for the pretrained text encoder. A weight decay of $1\times 10^{-4}$ was applied. A linear learning rate warmup was implemented for the first 5 epochs. Following the warmup, a Cosine Annealing with Warm Restarts [41] scheduler was used, with $T_{0}=10$ epochs and $T_{mult}=2$, and a minimum learning rate of $1\times 10^{-7}$. Gradients were clipped to a maximum L2 norm of 1.0. The loss function was the standard CLIP contrastive loss [42]. Mixed-precision training [43] was enabled. A full 25-epoch training run required approximately 2 days, with peak GPU memory consumption observed at roughly 14 GB per GPU.
4.2 Experimental setup
We evaluated GeoTreeCLIP against two advanced pretrained vision-language models: the original CLIP [13] and RemoteCLIP [24], a specialized VLM for remote sensing applications. To provide a fair comparison with models not inherently designed for time-series, we adopted an ensemble-like approach for CLIP and RemoteCLIP: features/probabilities were computed for each of the 12 monthly images independently, and the results were then averaged.
All models were evaluated on the GlobalGeoTree-10kEval benchmark using zero-shot and few-shot learning settings. Performance was measured using top-1 and top-5 prediction accuracy, with separate evaluations for each taxonomic level (family, genus, and species). To ensure robustness, we repeated each experiment 5 times using different random seeds and reported the mean accuracy and variance.
Zero-shot evaluation
For zero-shot evaluation, we assessed each model’s ability to classify samples at three taxonomic levels without specific training on the target categories [44].
Few-shot evaluation
For few-shot evaluation, we explored scenarios such as 1-shot learning, where the model is provided with only one labeled example per species. To implement this, we adopted a fine-tuning-based approach [45] using the pre-trained model. Specifically, we randomly sampled $k$ labeled examples per class (e.g., $k=1$, $k=3$) to form the support set and fine-tuned the visual encoder of the pre-trained model on this set. During fine-tuning, most of the visual encoder’s parameters were frozen, with only the last four transformer layers and the classification-related parameters remaining trainable. The text encoder was entirely frozen, leveraging the pre-trained textual embeddings for class labels. The fine-tuning was conducted for 10 epochs to balance adaptation and prevent overfitting due to the small support set. Afterward, the model was evaluated on the query set, which consisted of the remaining examples in the dataset.
For each query image, predictions were made by computing similarity scores between its visual embedding (extracted by the fine-tuned visual encoder) and the textual embeddings of the class labels. The class with the highest similarity score was assigned as the predicted label. Classification accuracy on the query set was then used to evaluate the model’s performance. This fine-tuning approach enables the model to adapt to the few-shot setting while retaining the benefits of the pre-trained representations.
This evaluation framework highlights the model’s ability to generalize effectively from limited labeled data, which is a crucial capability for real-world applications in biodiversity monitoring where obtaining large amounts of labeled data for every species is often infeasible.
4.3 Experimental results
4.3.1 Zero-shot evaluation
Table 3: Zero-shot evaluation on GlobalGeoTree-10kEval. Results are presented as mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP RemoteCLIP GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Family 10.80 $\pm$ 0.03 25.32 $\pm$ 0.05 1.11 $\pm$ 0.01 10.55 $\pm$ 0.04 20.99 $\pm$ 0.28 56.88 $\pm$ 0.42
Genus 1.09 $\pm$ 0.01 9.34 $\pm$ 0.01 1.11 $\pm$ 0.01 6.25 $\pm$ 0.02 18.39 $\pm$ 0.26 50.98 $\pm$ 0.41
Species 1.09 $\pm$ 0.01 7.02 $\pm$ 0.02 1.11 $\pm$ 0.01 6.25 $\pm$ 0.02 16.71 $\pm$ 0.25 47.52 $\pm$ 0.37
The results of zero-shot evaluation are presented in Table 3, clearly demonstrating the substantial improvements achieved by GeoTreeCLIP across all taxonomic levels. At the family level, GeoTreeCLIP achieves a top-1 accuracy of 20.99% and a top-5 accuracy of 56.88%. The performance gap is even more pronounced at the genus level. GeoTreeCLIP achieves a top-1 accuracy of 18.39% and a top-5 accuracy of 50.98%, outperforming CLIP (1.09% top-1, 9.34% top-5) and RemoteCLIP (1.11% top-1, 6.25% top-5) by a large margin. At the most challenging species level, GeoTreeCLIP still shows significant superiority, achieving a top-1 accuracy of 16.71% and a top-5 accuracy of 47.52%.
The experimental results reveal two critical patterns. First, the accuracy consistently declines as the taxonomic level becomes finer, reflecting the increasing complexity of distinguishing between closely related classes. This trend is evident across all models but is particularly severe for CLIP and RemoteCLIP, which exhibit minimal capability at the genus and species levels. In contrast, GeoTreeCLIP maintains stronger performance even at these fine-grained levels, likely due to its ability to learn and leverage the hierarchical relationships inherent in the taxonomic labels, as qualitatively supported by the feature embedding visualizations (see Appendix D and Figure 7). Second, the substantial performance gap between GeoTreeCLIP and the baseline models highlights the importance of domain-specific pretraining. GeoTreeCLIP’s ability to incorporate spatiotemporal and multispectral information enables it to address the unique challenges of tree species classification, which are beyond the scope of general-purpose models like CLIP and RemoteCLIP.
4.3.2 Few-shot evaluation
Table 4: Few-shot evaluation on GlobalGeoTree-10kEval. Results are presented as mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP RemoteCLIP GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
One-Shot Evaluation
Family 2.95 $\pm$ 0.01 15.06 $\pm$ 0.03 11.25 $\pm$ 0.01 23.85 $\pm$ 0.03 29.37 $\pm$ 0.07 69.38 $\pm$ 0.34
Genus 2.43 $\pm$ 0.01 8.14 $\pm$ 0.02 2.31 $\pm$ 0.01 7.68 $\pm$ 0.02 27.70 $\pm$ 0.11 64.40 $\pm$ 0.29
Species 1.67 $\pm$ 0.01 6.59 $\pm$ 0.03 1.94 $\pm$ 0.00 6.59 $\pm$ 0.02 25.80 $\pm$ 0.15 62.43 $\pm$ 0.25
Three-Shot Evaluation
Family 6.19 $\pm$ 0.01 23.50 $\pm$ 0.03 4.44 $\pm$ 0.02 17.02 $\pm$ 0.04 37.77 $\pm$ 0.23 75.49 $\pm$ 0.25
Genus 4.04 $\pm$ 0.01 12.76 $\pm$ 0.03 2.77 $\pm$ 0.03 11.70 $\pm$ 0.02 36.19 $\pm$ 0.22 72.50 $\pm$ 0.23
Species 3.41 $\pm$ 0.01 11.46 $\pm$ 0.03 1.88 $\pm$ 0.03 9.40 $\pm$ 0.06 33.67 $\pm$ 0.24 71.53 $\pm$ 0.23
The results in Table 4 demonstrate that providing even a small amount of labeled data for fine-tuning generally improves performance compared to the zero-shot setting across all models and taxonomic levels. GeoTreeCLIP consistently achieves the highest accuracies in both one-shot and three-shot scenarios. For instance, in the one-shot setting at the species level, GeoTreeCLIP reaches a top-1 accuracy of 25.80%, substantially outperforming CLIP (1.67%) and RemoteCLIP (1.94%). This advantage becomes even more pronounced with three shots, where GeoTreeCLIP’s species-level top-1 accuracy increases to 33.67%, while CLIP and RemoteCLIP show more modest gains to 3.41% and 1.88%, respectively.
Our few-shot experiments reveal distinct patterns across models when increasing from one to three shots. GeoTreeCLIP demonstrates substantial improvements across all taxonomic levels (species top-1 accuracy rising from 25.80% to 33.67%), while CLIP shows consistent but smaller gains. RemoteCLIP exhibits mixed results, including a slight decrease in family-level accuracy, suggesting difficulties effectively utilizing additional examples. Both baseline models demonstrate limited capacity to leverage few-shot supervision compared to GeoTreeCLIP, with only marginal improvements over their zero-shot performances (Table 3), particularly at finer taxonomic levels. This indicates that general pretraining approaches may not align sufficiently with the specific challenges of fine-grained tree species classification from remote sensing data, even when provided with in-domain examples.
The poor performance of CLIP and RemoteCLIP can likely be attributed to their design, which is optimized for RGB three-channel data and lacks the capability to process time-series information. Additionally, these models struggle with the small-patch classification tasks required for tree species identification. These limitations further emphasize the importance of introducing this benchmark for the global tree species classification task. More benchmark results on larger evaluation subsets can be found in Table 8 and Table 9 in Appendix C.
5 Ethics, Limitations and Impact
Ethics
GlobalGeoTree is constructed using publicly available data from sources like GBIF [46], Sentinel-2 [47], WorldClim [34], SoilGrids [33], and USGS SRTM [32], all under open licenses (CC0 1.0, CC-BY-4.0, CC-BY-NC-4.0). No personal or sensitive information is included, ensuring ethical use.
Limitations
Despite our efforts to ensure global coverage, the dataset exhibits geographic biases due to uneven species distribution data. Regions with longer histories of biodiversity documentation are overrepresented such as Europe and North America. The ambiguous boundary between trees and shrubs in botanical classification further complicates the dataset, as some samples may represent shrubs rather than trees. Tree species taxonomy is also subject to frequent revisions driven by new genetic evidence, which may misalign dataset labels with updated classifications over time. Although CLIP-based models can identify unseen species to some extent, such taxonomic shifts may still affect model interpretability and evaluation consistency. Additionally, the dataset relies on Sentinel-2 data from 2020, restricting its ability to capture long-term vegetation dynamics or recent disturbances. Future versions incorporating multi-year observations could better account for phenological changes and climate-driven impacts.
Potential impact
GlobalGeoTree holds great potential for advancing forest monitoring, biodiversity conservation, and climate change mitigation. By improving tree species mapping, it can support sustainable forest management, restoration planning, accurate carbon stock estimation, and biodiversity monitoring. However, there is a risk of misuse, such as in deforestation or logging. Responsible use and adherence to conservation principles are essential.
6 Availability and maintenance
The dataset access, pretrained model checkpoints, and all relevant codes are available in our github repository (https://github.com/MUYang99/GlobalGeoTree), which provides comprehensive tools for using these resources. The pretraining dataset GlobalGeoTree-6M and evaluation dataset GlobalGeoTree-10kEval are provided in WebDataset format [48] and hosted on Huggingface (https://huggingface.co/datasets/yann111/GlobalGeoTree). This format enables efficient online data streaming to train models without requiring full dataset downloads, facilitating large-scale machine learning workflows. It also integrates seamlessly with popular deep learning frameworks, improving accessibility and usability for researchers.
We are dedicated to maintaining and enhancing the dataset, addressing issues, and incorporating updates like new data sources or taxonomic revisions in future versions. The code and Huggingface repository will serve as the primary channels for updates and community feedback.
7 Conclusion and future work
In this paper, we introduced GlobalGeoTree, a large-scale, globally comprehensive dataset and benchmark for tree species classification. The dataset includes over 6 million geolocated tree occurrences spanning 21,001 species, paired with Sentinel-2 time series data and a rich set of auxiliary environmental variables. We also proposed GeoTreeCLIP, a baseline vision-language model specifically designed for this task, leveraging domain-specific pretraining on GlobalGeoTree-6M. Experimental results demonstrate that GeoTreeCLIP significantly outperforms existing advanced models in classification accuracy across all taxonomic levels, highlighting both the effectiveness of our approach and the importance of introducing this benchmark for global tree species classification.
Future work could explore several promising directions. Expanding the GlobalGeoTree with more recent data, additional satellite sensors (e.g., SAR data for structural information), and a broader range of auxiliary variables could enhance its utility. Investigating alternative vision-language model architectures, pretraining strategies, and methods for addressing the long-tail distribution could further improve classification accuracy, especially at the species level. Developing techniques for uncertainty estimation and improving model explainability are also critical areas for future work. Moreover, applying GlobalGeoTree and GeoTreeCLIP to real-world applications in biodiversity monitoring, conservation, and forest management could provide practical support and holds great potential.
Appendix
Appendix A GlobalGeoTree Dataset Statistics
The GlobalGeoTree dataset is a large-scale, multimodal resource for tree species classification. This section provides detailed statistics complementing the overview in the main paper.
A.1 Basic Statistics
The dataset encompasses a comprehensive collection of geolocated tree occurrences:
  • Total Samples: 6,263,345
  • Countries/Regions Covered: 221
  • Taxonomic Coverage: Families: 275, Genera: 2,734, Species: 21,001
A.2 Long-tail Distribution Analysis
The dataset exhibits a characteristic long-tail distribution across taxonomic levels as shown in Fig. 5. This highlights the challenge of classifying both common and rare taxa:
  • Family Level: The top 20% of families (55 families) account for 91.01% of all samples. Conversely, 24 families (8.73% of total families) have fewer than 10 samples each.
  • Genus Level: The top 20% of genera (546 genera) cover 96.65% of the samples. A significant portion, 760 genera (27.80% of total genera), have fewer than 10 samples.
  • Species Level: The distribution is most skewed at the species level, where the top 20% of species (4,200 species) comprise 97.21% of the samples. A majority of species, 11,611 species (55.29% of total species), have fewer than 10 samples.
This long-tail distribution underscores the importance of evaluation strategies, like those employed for GlobalGeoTree-10kEval, that explicitly consider species rarity.
Figure 5: Long-tail distribution across taxonomic levels in GlobalGeoTree.
A.3 Detailed Categorical Statistics
Below are statistics for key categorical attributes within the GlobalGeoTree dataset, illustrating geographical and taxonomic diversity and distribution.
Location (location)
  • Number of unique countries/regions: 221
  • Top 5 most frequent locations:
    1. 1.
      United States of America: 1,932,465
    2. 2.
      Australia: 506,179
    3. 3.
      Canada: 429,266
    4. 4.
      Colombia: 330,896
    5. 5.
      Russian Federation: 209,019
Functional Type (level0)
  • Number of unique functional types: 4
  • Distribution of functional types:
    • Deciduous Broadleaf: 3,582,456
    • Evergreen Broadleaf: 2,208,578
    • Evergreen Needleleaf: 447,568
    • Deciduous Needleleaf: 24,743
Taxonomic Family (level1_family)
  • Number of unique families: 275
  • Top 5 most frequent families:
    1. 1.
      Ericaceae: 423,365
    2. 2.
      Fabaceae: 384,038
    3. 3.
      Fagaceae: 362,317
    4. 4.
      Rosaceae: 355,950
    5. 5.
      Pinaceae: 320,415
Taxonomic Genus (level2_genus)
  • Number of unique genera: 2,734
  • Top 5 most frequent genera:
    1. 1.
      Cornus: 288,678
    2. 2.
      Quercus: 210,104
    3. 3.
      Pinus: 168,917
    4. 4.
      Vaccinium: 158,362
    5. 5.
      Prunus: 125,604
Taxonomic Species (level3_species)
  • Number of unique species: 21,001
  • Top 5 most frequent species:
    1. 1.
      Cornus acuminata: 180,120
    2. 2.
      Securidaca volubilis: 103,441
    3. 3.
      Cupania sylvatica: 99,797
    4. 4.
      Bourreria cumanensis: 96,304
    5. 5.
      Fagus sylvatica: 75,503
Appendix B Details of Evaluation Subsets
B.1 Overview and Construction
To enable robust benchmarking across various taxonomic diversity scales and species rarity, we constructed three evaluation subsets: GlobalGeoTree-10kEval, GlobalGeoTree-10kEval-300, and GlobalGeoTree-10kEval-900. These subsets were created by first categorizing all species in the GlobalGeoTree dataset into Rare, Common, and Frequent groups based on available sample counts (Section 3.3), then randomly selecting 30, 100, and 300 species per category, respectively. The primary GlobalGeoTree-10kEval benchmark (90 species) is featured in the main paper, while the larger subsets enable assessment of model scalability and performance on increasingly complex tasks. Detailed overviews of each subset’s composition are provided in Table 5, Table 6, and Table 7. The geographical distribution of the two additional evaluation sets is shown in Figure 6(a) and Figure 6(b).
B.2 GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
(a) GlobalGeoTree-10kEval-300
(b) GlobalGeoTree-10kEval-900
Figure 6: Geographic distributions of GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900.
Table 5: Overview of the GlobalGeoTree-10kEval evaluation subset. This subset comprises 30 species selected from each of the three rarity categories (Rare, Common, Frequent), totaling 90 unique species and 9,930 geolocated samples.
Category Example Species Num of Samples
Rare Acacia platycarpa 40
Adenanthos cuneatus 40
Adenocarpus decorticans 40
Rare (Total) 30 species selected 1,200
Common Abies religiosa 110
Aloe marlothii 110
Alternanthera sessilis 110
Common (Total) 30 species selected 3,300
Frequent Acer glabrum 181
Arctostaphylos glandulosa 181
Ardisia paniculata 181
Frequent (Total) 30 species selected 5,430
Total 90 species total 9,930
Table 6: Overview of the GlobalGeoTree-10kEval-300 evaluation subset. This subset comprises 100 species selected from each of the three rarity categories (Rare, Common, Frequent), totaling 300 unique species and 10,000 geolocated samples.
Category Example Species Num of Samples
Rare Abutilon wrightii 12
Acacia georgensis 12
Acacia loroloba 12
Rare (Total) 100 species selected 1,200
Common Acacia confusa 33
Acacia mucronata 33
Achyranthes spec 33
Common (Total) 100 species selected 3,300
Frequent Acacia dealbata 55
Acacia decurrens 55
Acacia polyphylla 55
Frequent (Total) 100 species selected 5,500
Total 300 species total 10,000
Table 7: Overview of the GlobalGeoTree-10kEval-900 evaluation subset. This subset comprises 300 species selected from each of the three rarity categories (Rare, Common, Frequent), totaling 900 unique species and 10,200 geolocated samples.
Category Example Species Num of Samples
Rare Abies hickelii 4
Abutilon auritum 4
Acacia adunca 4
Rare (Total) 300 species selected 1,200
Common Abies fraseri 11
Acacia echinula 11
Acacia falcata 11
Common (Total) 300 species selected 3,300
Frequent Abies amabilis 19
Abies balsamea 19
Acacia dealbata 19
Frequent (Total) 300 species selected 5,700
Total 900 species total 10,200
Appendix C Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
The evaluation results on the larger GlobalGeoTree-10kEval-300 (Table 8) and GlobalGeoTree-10kEval-900 (Table 9) subsets align with trends from the primary benchmark GlobalGeoTree-10kEval (Table 3 and 4). GeoTreeCLIP consistently outperforms CLIP and RemoteCLIP across all settings (zero-shot, one-shot, three-shot) and taxonomic levels (Family, Genus, Species). Despite lower absolute accuracies on these challenging subsets, especially GlobalGeoTree-10kEval-900, GeoTreeCLIP maintains a significant performance edge, highlighting the advantages of its domain-specific pretraining and tailored architecture for multimodal tree species classification.
Table 8: Zero-shot and Few-shot evaluation on GlobalGeoTree-10kEval-300. Results are presented as mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP RemoteCLIP GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Zero-Shot Evaluation
Family 7.07 $\pm$ 0.02 20.14 $\pm$ 0.03 1.35 $\pm$ 0.01 7.28 $\pm$ 0.08 12.55 $\pm$ 0.20 40.25 $\pm$ 0.30
Genus 2.34 $\pm$ 0.01 5.59 $\pm$ 0.03 0.57 $\pm$ 0.01 2.10 $\pm$ 0.05 9.26 $\pm$ 0.24 28.34 $\pm$ 0.25
Species 0.46 $\pm$ 0.00 2.06 $\pm$ 0.01 0.56 $\pm$ 0.01 1.70 $\pm$ 0.02 7.87 $\pm$ 0.20 25.20 $\pm$ 0.21
One-Shot Evaluation
Family 1.55 $\pm$ 0.00 6.80 $\pm$ 0.02 2.43 $\pm$ 0.01 10.30 $\pm$ 0.09 18.58 $\pm$ 0.21 50.41 $\pm$ 0.21
Genus 0.87 $\pm$ 0.01 3.59 $\pm$ 0.01 0.66 $\pm$ 0.01 3.91 $\pm$ 0.07 14.57 $\pm$ 0.26 41.92 $\pm$ 0.27
Species 0.63 $\pm$ 0.00 2.35 $\pm$ 0.01 0.61 $\pm$ 0.01 2.75 $\pm$ 0.03 13.31 $\pm$ 0.22 38.39 $\pm$ 0.26
Three-Shot Evaluation
Family 5.28 $\pm$ 0.01 16.12 $\pm$ 0.04 3.79 $\pm$ 0.01 11.49 $\pm$ 0.05 23.91 $\pm$ 0.26 57.97 $\pm$ 0.29
Genus 1.79 $\pm$ 0.01 6.02 $\pm$ 0.02 1.11 $\pm$ 0.02 3.60 $\pm$ 0.04 19.22 $\pm$ 0.28 50.35 $\pm$ 0.27
Species 1.33 $\pm$ 0.00 4.17 $\pm$ 0.01 0.74 $\pm$ 0.02 2.64 $\pm$ 0.02 17.90 $\pm$ 0.23 47.54 $\pm$ 0.29
Table 9: Zero-shot and Few-shot evaluation on GlobalGeoTree-10kEval-900. Results are presented as mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP RemoteCLIP GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Zero-Shot Evaluation
Family 4.77 $\pm$ 0.04 13.17 $\pm$ 0.05 1.10 $\pm$ 0.00 10.08 $\pm$ 0.02 7.62 $\pm$ 0.13 27.25 $\pm$ 0.05
Genus 0.69 $\pm$ 0.01 2.58 $\pm$ 0.02 0.04 $\pm$ 0.00 0.55 $\pm$ 0.02 4.36 $\pm$ 0.10 16.32 $\pm$ 0.11
Species 0.12 $\pm$ 0.00 0.59 $\pm$ 0.00 0.04 $\pm$ 0.00 0.51 $\pm$ 0.02 3.40 $\pm$ 0.09 11.83 $\pm$ 0.05
One-Shot Evaluation
Family 3.44 $\pm$ 0.02 9.76 $\pm$ 0.04 1.89 $\pm$ 0.04 4.69 $\pm$ 0.06 14.16 $\pm$ 0.26 41.81 $\pm$ 0.31
Genus 1.53 $\pm$ 0.01 2.82 $\pm$ 0.02 0.25 $\pm$ 0.01 1.21 $\pm$ 0.02 10.24 $\pm$ 0.21 31.45 $\pm$ 0.32
Species 0.45 $\pm$ 0.01 1.27 $\pm$ 0.01 0.22 $\pm$ 0.01 0.83 $\pm$ 0.01 8.18 $\pm$ 0.15 25.17 $\pm$ 0.28
Three-Shot Evaluation
Family 4.39 $\pm$ 0.02 12.93 $\pm$ 0.04 5.43 $\pm$ 0.09 13.29 $\pm$ 0.02 16.00 $\pm$ 0.14 46.07 $\pm$ 0.34
Genus 2.08 $\pm$ 0.01 5.25 $\pm$ 0.02 0.77 $\pm$ 0.02 4.22 $\pm$ 0.04 12.50 $\pm$ 0.21 37.42 $\pm$ 0.20
Species 1.37 $\pm$ 0.01 3.97 $\pm$ 0.01 0.34 $\pm$ 0.02 1.84 $\pm$ 0.06 10.23 $\pm$ 0.25 31.79 $\pm$ 0.17
Appendix D Qualitative Analysis: t-SNE Feature Embeddings Visualization
To qualitatively evaluate feature representations, we used t-SNE [49] to visualize image embeddings from samples in the GlobalGeoTree-10kEval-300 dataset, comparing GeoTreeCLIP with the original CLIP. A hierarchical analysis was conducted: first examining five randomly selected families, then five genera within Rosaceae, and finally five species within Prunus. Figure 7 shows GeoTreeCLIP (top row) consistently forming clearer, more distinct clusters compared to CLIP (bottom row). At the family level (left column), GeoTreeCLIP separates families like Hernandiaceae and Rosaceae better, while CLIP shows significant overlap. At the genus level (middle column), GeoTreeCLIP clusters genera such as Rhodotypos and Rubus distinctly, whereas CLIP merges them into a dense structure. At the species level (right column), GeoTreeCLIP demonstrates its ability to effectively distinguish between five Prunus species (P. avium, P. caroliniana, P. ilicifolia, P. laurocerasus, P. obtusata), whereas CLIP struggles to differentiate them. These results demonstrate that GeoTreeCLIP learns more semantically meaningful and discriminative representations across taxonomic levels, aligning with its superior quantitative performance.
Figure 7: t-SNE visualization of feature embeddings extracted by GeoTreeCLIP and Original CLIP across different taxonomic levels, using samples from GlobalGeoTree-10kEval-300.